# The dual nature of GHZ9: coexisting AGN and star formation activity in a remote X-ray source at z=10.145

### v3

@@ 6.1 The super-massive black hole in GHZ9
However, it has also been suggested that at high accretion rates, the hot X-ray emitting plasma undergoes a large photon supply from the accretion disk and the funnel walls. This would Compton-cool the plasma down to times lower values than in standard AGN coronae, resulting in reduced X-ray emission (Madau & Haardt, 2024), consistently with the widespread X-ray weakness of high redshift AGN discovered by JWST (Ananna et al., 2024; Maiolino et al., 2025; Mazzolari et al., 2024a). In this scenario, the X-ray bolometric corrections would be times larger than in standard quasars, leading to a bolometric luminosity for GHZ9 so high that it would again require a BH mass of to power it.
6.2 Accretion history and BH mass to stellar mass ratio
@@
If Eddington-limited, the BH mass estimated for GHZ9 would require an initial seed of 106 M⊙ at z=18 (yellow solid line from Fig. 4, see Valiante et al., 2016). Allowing for past super-Eddington accretion with fEdd = 1.5 (fEdd = 2) would alleviate the need to grow the black hole in GHZ9 from a 106 M⊙ heavy seed already in place at z=18, reducing the required masses to lighter seeds of 104 M⊙ (102 M⊙) at z = 25 (Valiante et al., 2016), as shown by the pink dashed (blue dot-dashed) line in Fig. 4. We note that Dayal (2024) and Huang et al. (2024) proposed a sub-Eddington accretion scenario onto a supermassive primordial black hole seed of 104 M⊙ for GHZ9.
@@
The high black hole mass points towards a high MBH/Mstar ratio (Fig. 4, right panel). When considering Mstar obtained with the best-fit fAGN=0.3, the MBH derived for = 2.3 ( = 1.8) implies an MBH/Mstar= 0.330.22 (0.100.07). Such a high ratio is in agreement with the analysis in Kovács et al. (2024), and consistent with typical values measured in high redshift AGN (e.g., Maiolino et al., 2024a; Furtak et al., 2024).
@@
We note that the MBH/Mstar ratio is affected by significant systematic uncertainties due to the assumptions made in deriving MBH, and to the wide range of Mstar values associated with the observed photometry at varying AGN contributions (Sect. 4). We first aim at deriving a stringent lower limit on MBH/Mstar. The maximal estimate of the stellar mass is obtained in the case fAGN=0 (Mstar = 7.2 108 M⊙). We stress that this scenario is conservative, as both the SED-fitting and the 4400Å continuum suggest a non-negligible AGN contribution to the SED. We assessed that this value is robust against additional systematics in the assumed IMF and SFH. In fact, the adoption of a top-heavy IMF, which is considered to be more appropriate for a low-metallicity object as GHZ9 (e.g., Chon et al., 2022; Trinca et al., 2024a), would decrease the stellar mass estimate. Unsurprisingly, the SFH is poorly constrained, but the fit obtained for GHZ9 does not appear to be significantly affected by the “outshining” effect which may lead to an underestimate of Mstar (Giménez-Arteaga et al., 2024). Indeed, it predicts 50% of the mass to have formed in a 400 Myr old burst, and the uncertainty considers the case with a fraction as low as 10% of Mstar forming in the ongoing burst. We then consider the most conservative MBH estimate, which is obtained under the assumptions of unabsorbed emission with fEdd=1, = 1.8 and including both the nominal uncertainty on the X-ray flux and a factor of 2 uncertainty on the bolometric correction (MBH = 4.8 107 M⊙). We find the stringent lower limit on MBH/Mstar is 0.07. The lower bound MBH/M2% (red star in Fig. 4) is significantly higher than expected from the Kormendy & Ho (2013) relation, and in line with the accelerated growth of black holes relative to stellar mass observed in high redshift AGN. Much higher ratios are obtained when relaxing the aforementioned assumptions, i.e. a steep spectral index = 2.3 leads to an MBH/Mstar ratio consistently 0.22, and as high as 0.48 when assuming the SED to be 60% (fAGN=0.6) contributed by AGN emission.
@@ 7 Conclusions
The inferences obtained from the X-ray and SED analysis suggest intriguing scenarios for the formation of SMBH and the coevolution with their host galaxies. Our stringent limit of MBH/M0.02 indicates an accelerated evolution of the BH mass compared to the stellar mass. However, without constraints on the Eddington ratio, the scenarios regarding the initial BH seed remain uncertain, as in the case for other high redshift AGN (Larson et al., 2023b; Maiolino et al., 2024a; Bogdán et al., 2024).
@@
We also note that GHZ9 belongs to a well defined, homogeneous photometric sample selected based on pure Lyman-Break color selection criteria (Castellano et al., 2022, 2023). As described in Napolitano et al. (2025), program GO-3073 confirmed all the six objects from the parent GLASS-JWST sample at , plus two additional sources from alternative photometric selections. Out of these objects, at least GHZ9 is a confirmed AGN. Despite the low–number statistics, this is in line with the growing consensus that about 10-15% of bright, high redshift galaxies host AGNs (e.g. Maiolino et al., 2024b).
@@
It is also tempting to place these findings in the context of other high redshift galaxy observations. The most obvious analog is GHZ2, another bright object at , which displays strong UV emission lines (Castellano et al., 2024). As shown in Fig. 3, GHZ2 and GHZ9 are essentially indistinguishable in terms of UV line ratios and EW - the only notable difference being the He II line, with an EW18Å in GHZ9 which is significantly higher than in GHZ2 (EW5Å). In both cases, these objects are compact and N–enhanced, with a C/O ratio in line with expectations for low-metallicity galaxies at high redshift. Unfortunately, lacking an X-ray analysis and high-resolution JWST spectra, it is not possible to ascertain whether GHZ2 also hosts an AGN - future observations are clearly needed to investigate this.
@@
GHZ9 also shares common properties with the bright object GNz11 at z=10.6, which is compact and N-enhanced, albeit with a less extreme UV spectrum (Bunker et al., 2023), and has broad line components and extreme gas densities, indicating the presence of a BLAGN (Maiolino et al., 2024a).
@@
Based on their UV line emission properties alone, GHZ9 (and GHZ2) are similar to the class of strong C IV-emitters at lower redshifts (Izotov et al., 2024; Topping et al., 2024a, b). There is no evidence that these objects host an AGN, and their line ratios seem to indicate that they experience a dense starburst. This leaves open the possibility that many, if not all, of these newly-discovered sources also host AGN, but their active nuclei remain elusive due to X-ray weakness, beamed, or absorbed emission (e.g., Madau et al., 2024; Maiolino et al., 2024b). A comprehensive scenario connecting GHZ9, other strong C IV emitters like GHZ2, and the broader category of N-enhanced objects, including GNz11 and others (Isobe et al., 2023; Schaerer et al., 2024; Topping et al., 2024b) is currently missing. However, a composite nature involving both AGN activity and star formation in a low-metallicity, dense ISM enriched by massive or super-massive stars appears consistent with the available evidence (Charbonnel et al., 2023; D’Antona et al., 2023; Marques-Chaves et al., 2024; Calabrò et al., 2024; Zavala et al., 2025).
@@
These scenarios deserve an in-depth analysis with future observations. In this context, GHZ9, which is the only z9 object showing both a highly-ionizing, N-enhanced spectrum and X-ray emission, is the ideal target to improve constraints on SMBH-host coevolution and SMBH seeding mechanisms. Specifically, we can individuate two main directions for progress: 1) constraining the physical properties of the host, and the contribution of the AGN to the total UV/optical emission; 2) obtaining independent constraints on MBH and, thus, the Eddington ratio.
@@
The physical conditions of the star-forming ISM (density, temperature, abundances) can be assessed through high-resolution spectroscopy, to accurately constrain the origin of the copious amounts of ionizing photons (e.g., Ji et al., 2024). Medium or high resolution NIRSpec spectroscopy, as well as MIRI MRS, can search for broad-line components in the permitted lines of GHZ9 and other similar sources to assess the AGN fraction, currently the main uncertainty in deriving MBH/Mstar. Most importantly, estimating MBH from broad lines would allow determination of the Eddington ratio in combination with the X-ray luminosity (e.g., Lusso et al., 2010), providing a direct constraint on the seeding mechanisms of the SMBH in GHZ9. Finally, ALMA can measure the dynamical mass using FIR lines, enabling direct comparison between MBH/Mstar and MBH/Mdyn, which would constrain the relative timescales of BH and stellar mass growth (e.g., Pensabene et al., 2020).
@@
Significant progress in determining the X-ray properties of GHZ9 and similar objects (e.g., UHZ1 Goulding et al., 2023) will have to await next-generation X-ray imaging satellites with large collecting areas and arcsec angular resolution, such as the Advanced X-ray Imaging Satellite (AXIS), a probe-class mission currently under evaluation at NASA (Marchesi et al., 2020; Reynolds et al., 2023). For the time being, as in the case of GHZ9, a detailed analysis of JWST-selected candidates discovered in fields with deep Chandra imaging will be essential to further constrain different scenarios on the early stages of galaxy-AGN coevolution.

### v4

@@ 6.1 The Supermassive Black Hole in GHZ9
However, it has also been suggested that at high accretion rates, the hot X-ray emitting plasma undergoes a large photon supply from the accretion disk and the funnel walls. This would Compton-cool the plasma down to times lower values than in standard AGN coronae, resulting in reduced X-ray emission (Madau & Haardt, 2024), consistent with the widespread X-ray weakness of high-redshift AGN discovered by JWST (Ananna et al., 2024; Mazzolari et al., 2024a; Maiolino et al., 2025). In this scenario, the X-ray bolometric corrections would be times larger than in standard quasars, leading to a bolometric luminosity for GHZ9 that is so high that it would again require a BH mass of to power it.
6.2 Accretion History and BH-mass-to-stellar-mass Ratio
@@
If Eddington-limited, the BH mass estimated for GHZ9 would require an initial seed of 106 M⊙ at z=18 (yellow solid line from Figure 4, see Valiante et al., 2016). Allowing for past super-Eddington accretion with fEdd = 1.5 (fEdd = 2) would alleviate the need to grow the BH in GHZ9 from a 106 M⊙ heavy seed already in place at z=18, reducing the required masses to lighter seeds of 104 M⊙ (102 M⊙) at z = 25 (Valiante et al., 2016), as shown by the pink dashed (blue dot-dashed) line in Figure 4. We note that Dayal (2024) and Huang et al. (2024) proposed a sub-Eddington accretion scenario onto a supermassive primordial BH seed of 104 M⊙ for GHZ9.
@@
The high BH mass points towards a high MBH/Mstar ratio (Figure 4, right panel). When considering Mstar obtained with the best-fit fAGN=0.3, the MBH derived for = 2.3 ( = 1.8) implies an MBH/Mstar= 0.330.22 (0.100.07). Such a high ratio is in agreement with the analysis in Kovács et al. (2024), and consistent with typical values measured in high-redshift AGN (e.g., Furtak et al., 2024; Maiolino et al., 2024a).
@@
We note that the MBH/Mstar ratio is affected by significant systematic uncertainties due to the assumptions made in deriving MBH, and to the wide range of Mstar values associated with the observed photometry at varying AGN contributions (Section 4). We first aim at deriving a stringent lower limit on MBH/Mstar. The maximal estimate of the stellar mass is obtained in the case fAGN=0 (Mstar = 7.2 108 M⊙). We stress that this scenario is conservative because both the SED fitting and the 4400 Å continuum suggest a non-negligible AGN contribution to the SED. We assessed that this value is robust against additional systematics in the assumed IMF and SFH. In fact, the adoption of a top-heavy IMF, which is considered to be more appropriate for a low-metallicity object as GHZ9 (e.g., Chon et al., 2022; Trinca et al., 2024a), would decrease the stellar mass estimate. Unsurprisingly, the SFH is poorly constrained, but the fit obtained for GHZ9 does not appear to be significantly affected by the “outshining” effect, which may lead to an underestimate of Mstar (Giménez-Arteaga et al., 2024). Indeed, it predicts 50% of the mass to have formed in a 400 Myr old burst, and the uncertainty considers the case with a fraction as low as 10% of Mstar forming in the ongoing burst. We then consider the most conservative MBH estimate, which is obtained under the assumptions of unabsorbed emission with fEdd=1, = 1.8 and including both the nominal uncertainty on the X-ray flux and a factor of 2 uncertainty on the bolometric correction (MBH = 4.8 107 M⊙). We find the stringent lower limit on MBH/Mstar is 0.07. The lower bound MBH/M2% (red star in Figure 4) is significantly higher than expected from the Kormendy & Ho (2013) relation, and in line with the accelerated growth of BHs relative to stellar mass observed in high-redshift AGN. Much higher ratios are obtained when relaxing the aforementioned assumptions, i.e., a steep spectral index = 2.3 leads to an MBH/Mstar ratio consistently 0.22, and as high as 0.48 when assuming the SED to be 60% (fAGN=0.6) contributed by AGN emission.
@@ 7 Conclusions
The inferences obtained from the X-ray and SED analysis suggest intriguing scenarios for the formation of SMBH and the coevolution with their host galaxies. Our stringent limit of MBH/M0.02 indicates an accelerated evolution of the BH mass compared to the stellar mass. However, without constraints on the Eddington ratio, the scenarios regarding the initial BH seed remain uncertain, as in the case for other high-redshift AGN (Larson et al., 2023b; Bogdán et al., 2024; Maiolino et al., 2024a).
@@
We also note that GHZ9 belongs to a well-defined, homogeneous photometric sample selected based on pure Lyman-Break color selection criteria (Castellano et al., 2022, 2023). As described in Napolitano et al. (2025), program GO-3073 confirmed all the six objects from the parent GLASS-JWST sample at , plus two additional sources from alternative photometric selections. Out of these objects, at least GHZ9 is a confirmed AGN. Despite the low–number statistics, this is in line with the growing consensus that about 10-15% of bright, high-redshift galaxies host AGNs (e.g. Maiolino et al., 2024b).
@@
It is also tempting to place these findings in the context of other high-redshift galaxy observations. The most obvious analog is GHZ2, another bright object at , which displays strong UV emission lines (Castellano et al., 2024). As shown in Figure 3, GHZ2 and GHZ9 are essentially indistinguishable in terms of UV line ratios and EW - the only notable difference being the He II line, with an EW18 Å in GHZ9 which is significantly higher than in GHZ2 (EW5 Å). In both cases, these objects are compact and N–enhanced, with a C/O ratio in line with expectations for low-metallicity galaxies at high redshift. Unfortunately, lacking an X-ray analysis and high-resolution JWST spectra, it is not possible to ascertain whether GHZ2 also hosts an AGN - future observations are clearly needed to investigate this.
@@
GHZ9 also shares common properties with the bright object GN-z11 at z=10.6, which is compact and N-enhanced, albeit with a less extreme UV spectrum (Bunker et al., 2023), and has broad-line components and extreme gas densities, indicating the presence of a BLAGN (Maiolino et al., 2024a).
@@
Based on their UV line-emission properties alone, GHZ9 (and GHZ2) are similar to the class of strong C IV emitters at lower redshifts (Izotov et al., 2024; Topping et al., 2024a, b). There is no evidence that these objects host an AGN, and their line ratios seem to indicate that they experience a dense starburst. This leaves open the possibility that many, if not all, of these newly-discovered sources also host AGN, but their active nuclei remain elusive due to X-ray weakness, beamed, or absorbed emission (e.g., Madau et al., 2024; Maiolino et al., 2024b). A comprehensive scenario connecting GHZ9, other strong C IV emitters such as GHZ2, and the broader category of N-enhanced objects, including GN-z11 and others (Isobe et al., 2023; Schaerer et al., 2024; Topping et al., 2024b) is currently missing. However, a composite nature involving both AGN activity and star formation in a low-metallicity, dense interstellar medium (ISM) enriched by massive or supermassive stars appears consistent with the available evidence (Charbonnel et al., 2023; D’Antona et al., 2023; Boorman et al., 2024; Calabrò et al., 2024; Marques-Chaves et al., 2024; Zavala et al., 2025).
@@
These scenarios deserve an in-depth analysis with future observations. In this context, GHZ9, which is the only z9 object showing both a highly-ionizing, N-enhanced spectrum and X-ray emission, is the ideal target to improve constraints on SMBH-host coevolution and SMBH seeding mechanisms. Specifically, we can individuate two main directions for progress: (1) constraining the physical properties of the host, and the contribution of the AGN to the total UV/optical emission; and (2) obtaining independent constraints on MBH, and thus the Eddington ratio.
@@
The physical conditions of the star-forming ISM (i.e., density, temperature, abundances) can be assessed through high-resolution spectroscopy, to accurately constrain the origin of the copious amounts of ionizing photons (e.g., Ji et al., 2024). Medium or high-resolution NIRSpec spectroscopy, as well as MIRI Medium Resolution Spectroscopy, can search for broad-line components in the permitted lines of GHZ9 and other similar sources to assess the AGN fraction, currently the main uncertainty in deriving MBH/Mstar. Most importantly, estimating MBH from broad lines would allow determination of the Eddington ratio in combination with the X-ray luminosity (e.g., Lusso et al., 2010), providing a direct constraint on the seeding mechanisms of the SMBH in GHZ9. Finally, ALMA can measure the dynamical mass using FIR lines, enabling direct comparison between MBH/Mstar and MBH/Mdyn, which would constrain the relative timescales of BH and stellar-mass growth (e.g., Pensabene et al., 2020).
@@
Significant progress in determining the X-ray properties of GHZ9 and similar objects (e.g., UHZ1 Goulding et al., 2023) will have to await next-generation X-ray imaging satellites with large collecting areas and arcsec angular resolution, such as the Advanced X-ray Imaging Satellite, a probe-class mission currently under evaluation at NASA (Marchesi et al., 2020; Reynolds et al., 2023). For the time being, as in the case of GHZ9, a detailed analysis of JWST-selected candidates discovered in fields with deep Chandra imaging will be essential to further constrain different scenarios on the early stages of galaxy-AGN coevolution.